# Simplicial complexes with many facets are vertex decomposable

### v1

@@ Abstract.
Let $\Delta$ be a pure simplicial complex on $n$ vertices having dimension $d$ and codimension $c=n-d-1$ in the simplex. Terai and Yoshida proved that if the number of facets of $\Delta$ is at least $\binom{n}{c}-2c+1$, then $\Delta$ is Cohen-Macaulay. We improve this result by showing that these hypotheses imply the stronger condition that $\Delta$ is vertex decomposable. We give examples to show that this bound is optimal, and that the conclusion cannot be strengthened to the class of matroids or shifted complexes. We explore an application to Simon’s Conjecture and discuss connections to other results from the literature.
@@ 1. Introduction
Suppose $\Delta$ is a simplicial complex on vertex set $[n]=\{1,\ldots,n\}$. Recall that $\Delta$ is pure if all the facets of $\Delta$ (the maximal elements of $\Delta$ under inclusion) have the same cardinality. As first shown by Terai and Yoshida [11, Theorem 3.1], if the number of facets of a pure simplicial complex $\Delta$ is “large enough,” then $\Delta$ is Cohen-Macaulay. In other words, the simplicial complex $\Delta$ enjoys additional topological properties that can be detected by simply counting the number of facets.
In this short note we strengthen Terai and Yoshida’s results. As in their paper, we express our theorem in terms of the codimension of $\Delta$ relative to the simplex on the same vertex set. Recall that the dimension of $\Delta$ is given by $\dim(\Delta)=\max\{\dim(F)\leavevmode\nobreak\ |\leavevmode\nobreak\ F\in\Delta\}$, where $\dim(F)=|F|-1$. If $\Delta$ has $n$ vertices, then the codimension of $\Delta$, denoted $\operatorname{{codim}}{\Delta}$, is given by $c=n-\dim(\Delta)-1$ (the number of vertices in the complement of a facet). Our main result can now be stated as follows.
@@
Vertex decomposable simplicial complexes, which were introduced by Provan and Billera in [9], are a recursively defined family of simplicial complexes (see Section 2 for the formal definition). Terai and Yoshida [11, Theorem 3.1] showed that, under the same hypotheses as Theorem 1.1, the simplicial complex $\Delta$ is Cohen-Macaulay. Their result was stated in terms of the multiplicity of the Stanley-Reisner ring, which agrees with the number of facets of $\Delta$ when $\Delta$ is pure. Since being vertex decomposable implies being Cohen-Macaulay, our main theorem strengthens their result. To prove [11, Theorem 3.1], Terai and Yoshida used an algebraic approach, employing the Stanley-Reisner correspondence between simplicial complexes and square-free monomials. Our proof, in contrast, requires only simple tools from topological combinatorics.
The bound in Theorem 1.1 is tight. In Example 4.1 we describe a family of simplicial complexes on $n$ vertices of codimension $c$ and with $\binom{n}{c}-2c$ facets, each of which fails to be vertex decomposable. In Remark 4.3 we also show that Theorem 1.1 cannot be strengthened to conclude that $\Delta$ is a matroid complex or shifted complex (both of which classes are known to be vertex decomposable), even if we delete only two facets.
It is not hard to see that a simplicial complex $\Delta$ on vertex set $[n]$ with dimension $d$ and codimension $c$ can have more than $\binom{n}{c}-2c$ faces of dimension $d$, but fail to be pure. For instance, consider the 1-dimensional complex consisting of a triangle and an isolated vertex. However, we observe in Lemma 4.4 that if we require at least $\binom{n}{c}-c$ facets of this top dimension, then purity is guaranteed. This leads to Corollary 4.5 for arbitrary simplicial complexes, which strengthens [11, Theorem 2.1].
@@
Our paper is organized as follows. In Section 2 we recall the necessary definitions and background, in particular relevant properties of vertex decomposable simplicial complexes. In Section 3 we prove Theorem 1.1. In Section 4, we address the optimality of Theorem 1.1, interpret our result in the language of square-free monomial ideals, and discuss the connection to Simon’s Conjecture and Lasoń’s result.
@@ 2. Background
In this section, we recall additional relevant background needed to prove our main result. We refer to [6] or [12] for additional background on combinatorial commutative algebra, and to [1] for topological combinatorics. We continue to use the notation and definitions from the introduction.
Let $\Delta$ be a simplicial complex on vertex set $V$, which we typically take as $V=[n]:=\{1,2,\dots,n\}$. An element $F\in\Delta$ is called a face, and a face not properly contained in any other face is called a facet. For each $i\in[n]$ we always assume that $\{i\}$ is a face of $\Delta$, which we call a vertex and denote simply as $i$. If $F_{1},\ldots,F_{s}$ is a complete list of the facets of $\Delta$, then we use the notation $\Delta=\langle F_{1},\ldots,F_{s}\rangle$. After fixing a vertex $x\in[n]$, we can define two other simplicial complexes from $\Delta$. Namely, the link of $\{x\}$ is the simplicial complex
${\rm link}_{\Delta}(\{x\})=\{G\in\Delta\leavevmode\nobreak\ |\leavevmode\nobreak\ x\notin G,\leavevmode\nobreak\ \leavevmode\nobreak\ G\cup\{x\}\in\Delta\},$
@@
${\rm del}_{\Delta}(\{x\})=\{G\in\Delta\leavevmode\nobreak\ |\leavevmode\nobreak\ x\notin G\}.$
As above we will abuse notation and write ${\rm link}_{\Delta}(x)$ (respectively, ${\rm del}_{\Delta}(x)$) instead of ${\rm link}_{\Delta}(\{x\})$ (respectively, ${\rm del}_{\Delta}(\{x\})$). A vertex $x$ is a shedding vertex if for any facet $F\in\Delta$ with $x\in F$, there exists a vertex $y\not\in F$ such that $(F\setminus\{x\})\cup\{y\}$ is also a face of $\Delta$. Equivalently, $x$ is a shedding vertex if every facet of ${\rm del}_{\Delta}(x)$ is a facet of $\Delta$.
@@
    $\Delta=\{\emptyset\}$, or $\Delta$ is a simplex, i.e., $\Delta=\langle F\rangle$ for some $F\subseteq[n]$, or
@@
We next collect some results regarding vertex decomposable complexes that will be useful for our study. Recall that a simplicial complex $\Delta$ is a cone if there exists a vertex $v\in V$ such that $v\in F$ for every facet $F\in\Delta$. In this case there exists a simplicial complex $\Gamma$ on vertex set $V\backslash\{v\}$ such that every face of $\Delta$ is either a face of $\Gamma$, or else of the form $F\cup\{v\}$ for some face $F\in\Gamma$. In this case we write $\Delta=\Gamma\ast v$. In the statement below, for $k\leq\dim(\Delta)$ the $k$-skeleton of a pure simplicial complex $\Delta$ is the simplicial complex whose facets are the faces of $\Delta$ having dimension $k$.
@@
    [9, Proposition 3.1.1] All 0-dimensional complexes are vertex decomposable.
@@
    [9, Proposition 3.1.2] A pure $1$-dimensional complex $\Delta$ is vertex decomposable if and only if $\Delta$ is connected.
@@
    [13, Lemma 3.10] Any $k$-skeleton of a vertex decomposable complex $\Delta$ is vertex decomposable.
@@
    [9, Proposition 2.4] A cone $\Delta=\Gamma\ast v$ is vertex decomposable if and only if $\Gamma$ is vertex decomposable.
@@
The following lemma is immediate from the recursive definition of vertex-decomposability, and we will use it without comment.
@@
If $\mathcal{F}$ is a family of simplicial complexes, so that every $\Delta\in\mathcal{F}$ is either a simplex or has a shedding vertex v with the property that both $\operatorname{link}_{\Delta}(v)$ and $\operatorname{del}_{\Delta}(v)$ are in $\mathcal{F}$, then every complex in $\mathcal{F}$ is vertex-decomposable.
@@ 3. Proof of Main Result
In this section we provide the proof of Theorem 1.1. In what follows, we let $\Delta$ be a pure simplicial complex on $n$ vertices with codimension $c=n-\dim\Delta-1$. If $F$ is a set of $\dim(\Delta)+1$ vertices that is not a face of $\Delta$, we say that $F$ is an antifacet. Write $e(\Delta)$ for the number of facets of $\Delta$ and $\bar{e}(\Delta)$ for the number of antifacets of $\Delta$. Note that $\bar{e}(\Delta)=\binom{n}{c}-e(\Delta)$.
@@
If $n\geq c+2>3$ and $\bar{e}(\Delta)\leq 2c-1$, then $\Delta$ has at least two vertices that each avoids at most $32c-$ 3 antifacets.
@@
Counting by antifacets, we see that $\left|\mathcal{P}\right|\leq c\cdot(2c-1)=2c^{2}-c$.
Suppose for contradiction that every vertex except for some vertex $w$ avoids more than $32c-$ 3 antifacets. Then
@@
As $22c^{2}-c\geq 2c^{2}-$ 2 holds only if $c\leq 2$, we have reduced to the case $c=2$. But in this case, the left-most inequality of (3.1) gives us that $2\cdot 3\geq(n-1)\cdot 2$, so we are left with the case $n=4$.
Hence we are left with the case that $\Delta$ is a one-dimensional complex on $4$ vertices. In this case, we have $\bar{e}(\Delta)\leq 3$ and $c=2$, and so we need to show there must be at least two vertices that each avoids at most $12c-3=$ 1 antifacets. Viewing the antifacets as edges in a graph on four vertices, the antifacets give a graph on four vertices where every vertex has degree at most two (since $\Delta$ cannot have an isolated vertex). We finish the proof with the following easily-verified observation: For any graph on four vertices with three or fewer edges, where the degree of each vertex is $\leq 2$, there are at least two vertices that are each disjoint from at most one edge. ∎
@@
We make some other easy observations. Recall that a minimal nonface of $\Delta$ is a nonface of $\Delta$ that is minimal under inclusion with respect to this property. Note that if $\{v,w\}$ is a minimal nonface of $\Delta$ of cardinality two, then $w$ is not a vertex of $\operatorname{link}_{\Delta}(v)$.
The following lemma says that minimal nonfaces of cardinality two are rare when $\dim(\Delta)\neq 1$ and the number of antifacets is small enough (equivalently, if the number of facets is large enough relative to $n$).
@@
If $\bar{e}(\Delta)\leq 2c-1$, then no minimal nonface of $\Delta$ has fewer than $\dim(\Delta)$ vertices, and at most one minimal nonface has $\dim(\Delta)$ vertices.
@@
A nonface of size $\dim(\Delta)-1$ is contained in $\binom{c+2}{2}$ antifacets. But $(c+2)(c+1)/2$ is greater than $12c-$ 1 for all $c$.
Each minimal nonface with $\dim(\Delta)$ vertices is contained in $c+1$ antifacets, and a pair of such minimal nonfaces is contained in at most one common antifacet. But $12(c+1)-$ 1 is greater than $12c-$ 1, contradicting our assumptions. The result follows. ∎
@@
Suppose $\dim(\Delta)\geq 1$, and let $v$ be a vertex of $\Delta$. If $\bar{e}(\Delta)\leq 2c-1$ and $b$ is the codimension of $\operatorname{link}_{\Delta}(v)$, then $\bar{e}(\operatorname{link}_{\Delta}(v))\leq 2b-1$.
@@
Note that since $\Delta$ is pure, we have that $\dim(\operatorname{link}_{\Delta}(v))=\dim(\Delta)-1$. Also note that the statement is trivial if $\dim(\Delta)\leq 1$, since in this case $\bar{e}(\operatorname{link}_{\Delta}(v))=0$. If $\dim(\Delta)>2$, then by Lemma 3.3 we have that $\Delta$ has no missing edge, so that $b=(n-1)-(\dim(\Delta)-1)-1=c$. The result then follows by noting that if $F$ is an antifacet of $\operatorname{link}_{\Delta}(v)$, then $F\cup\{v\}$ is an antifacet of $\Delta$. If $\dim(\Delta)=2$, then again by Lemma 3.3 we have that $b=c-1$ (so also that $c\geq 1$), and there is a unique minimal nonfacet $\{v,w\}$ having two vertices. Now, as before there are $\bar{e}(\operatorname{link}_{\Delta}(v))$ antifacets of $\Delta$ containing $v$ but not $w$, and an additional $n-2$ antifacets of $\Delta$ containing $\{v,w\}$. We obtain that $\bar{e}(\operatorname{link}_{\Delta}(v))\leq\bar{e}(\Delta)-(n-2)\leq(2c-1)-(c+1)=2b-c$, which yields the desired. ∎
@@
We collect one more preliminary result.
@@
    Any vertex that is contained in at least $c$ antifacets is a shedding vertex;
@@
Recall that $\Delta$ is assumed to be a pure $d$-dimensional simplicial complex on vertex set $[n]$ and codimension $c$, with at least $\binom{n}{c}-2c+1$ facets, so that $\bar{e}(\Delta)\leq 2c-1$. The statement can be checked directly for all simplicial complexes on $n\leq 3$ vertices. If $c=0$ or $1$, then $\Delta$ is a simplex, or a simplex boundary with at most one face removed, and the result follows by Lemma 2.4. On the other hand, if $c=n-1$, then $\dim(\Delta)=0$, and the result is immediate.
It is also convenient to handle the case where $\dim(\Delta)=1$ separately. By Lemma 2.3, a pure $1$-dimensional complex is vertex decomposable if and only if it is connected. Supppose $\Delta$ is disconnected, where the smallest connected component $\Gamma$ has $k$ vertices. As $\Delta$ is pure, we have $k\geq 2$. Now there is no edge between $\Gamma$ and its vertex complement, so $\Delta$ has at least $k\cdot(n-k)\geq 2\cdot(n-2)=2c$ antifacets, and we see that $\Delta$ fails the hypotheses of the theorem.
Moving forward, we can assume $\dim(\Delta)\geq 2$ and that $3<c+2\leq n$. To prove the theorem we wish to find a shedding vertex $v$ so that $\operatorname{link}_{\Delta}(v)$ and $\operatorname{del}_{\Delta}(v)$ also satisfy the hypotheses of the theorem. Note that we are using Lemma 2.5. First observe that for any vertex $v$, ${\rm link}_{\Delta}(v)$ satisfies the hypotheses by Corollary 3.4.
We next turn to the deletion. For this note that if $v$ is a shedding vertex, then by definition $\dim(\operatorname{del}_{\Delta}(v))=\dim(\Delta)$. Since the number of vertices of $\operatorname{del}_{\Delta}(v)$ is $n-1$, we have that the codimension of $\operatorname{del}_{\Delta}(v)$ is $c-1$. Hence to complete the proof we need to find a shedding vertex $v$ satisfying $\bar{e}(\operatorname{del}_{\Delta}(v))\leq 2c-3$.
If every (or even all but one) vertex of $\Delta$ is a shedding vertex, then we are done by Lemma 3.1. Otherwise, Lemma 3.6 yields a shedding vertex $v$ that is contained in at least $c$ antifacets, so that $\bar{e}(\operatorname{del}_{\Delta}(v))\leq 2c-1-c$. In either case, we have produced the desired shedding vertex. ∎

### v2

@@ Abstract.
Suppose $\Delta$ is a pure simplicial complex on $n$ vertices having dimension $d$ and let $c=n-d-1$ be its codimension in the simplex. Terai and Yoshida proved that if the number of facets of $\Delta$ is at least $\binom{n}{c}-2c+1$, then $\Delta$ is Cohen-Macaulay. We improve this result by showing that these hypotheses imply the stronger condition that $\Delta$ is vertex decomposable. We give examples to show that this bound is optimal, and that the conclusion cannot be strengthened to the class of matroids or shifted complexes. We explore an application to Simon’s Conjecture and discuss connections to other results from the literature.
@@ 1. Introduction
Suppose $\Delta$ is a simplicial complex on vertex set $[n]=\{1,\ldots,n\}$. Recall that $\Delta$ is pure if all the facets of $\Delta$ (the maximal elements of $\Delta$ under inclusion) have the same cardinality. As first shown by Terai and Yoshida [14, Theorem 3.1], if the number of facets of $\Delta$ is “large enough,” then $\Delta$ is Cohen-Macaulay. In other words, the simplicial complex $\Delta$ enjoys additional topological properties that can be detected by simply counting the number of facets.
In this short note we strengthen Terai and Yoshida’s results. As in their paper, we express our theorem in terms of the codimension of $\Delta$ relative to the simplex on the same vertex set. Recall that the dimension of $\Delta$ is given by $\dim\Delta=\max\{\dim F~{}|~{}F\in\Delta\}$, where $\dim F=|F|-1$. If $\Delta$ has $n$ vertices, then the codimension of $\Delta$, denoted $\operatorname{{codim}}{\Delta}$, is given by $c=n-\dim\Delta-1$ (the number of vertices in the complement of a facet). Our main result can now be stated as follows.
@@
The class of vertex decomposable simplicial complexes was introduced by Provan and Billera in [12], and have a recursive definition (see Section 2 for the formal statement). Terai and Yoshida [14, Theorem 3.1] showed that, under the same hypotheses as Theorem 1.1, the simplicial complex $\Delta$ is Cohen-Macaulay. Their result was stated in terms of the multiplicity of the Stanley-Reisner ring, which agrees with the number of facets of $\Delta$ when $\Delta$ is pure. Since being vertex decomposable implies being Cohen-Macaulay, our main theorem strengthens their result. To prove [14, Theorem 3.1], Terai and Yoshida used an algebraic approach, employing the Stanley-Reisner correspondence between simplicial complexes and square-free monomials. Our proof, in contrast, requires only simple tools from topological combinatorics.
The bound in Theorem 1.1 is tight. In Example 4.1 we describe a family of simplicial complexes on $n$ vertices of codimension $c$ with $\binom{n}{c}-2c$ facets, each of which fails to be vertex decomposable. In Remark 4.3 we also show that Theorem 1.1 cannot be strengthened to conclude that $\Delta$ is a matroid complex or shifted complex (both of which are classes known to be vertex decomposable), even if we require $\binom{n}{c}-2$ facets.
It is not hard to see that a simplicial complex $\Delta$ on vertex set $[n]$ with dimension $d$ and codimension $c$ can have more than $\binom{n}{c}-2c$ faces of dimension $d$, but fail to be pure. For instance, consider the $1$-dimensional complex consisting of a triangle and an isolated vertex. However, we observe in Lemma 4.4 that if we require at least $\binom{n}{c}-c$ facets of this top dimension, then purity is guaranteed. This leads to Corollary 4.5 for arbitrary simplicial complexes, which strengthens [14, Theorem 2.1].
@@
Our paper is organized as follows. In Section 2 we provide the necessary definitions and background, and in particular recall relevant properties of vertex decomposable simplicial complexes. In Section 3 we prove Theorem 1.1. In Section 4, we address the optimality of Theorem 1.1, interpret our result in the language of square-free monomial ideals, and discuss the connection to Simon’s Conjecture and Lasoń’s result.
@@ 2. Background
In this section, we recall additional relevant definitions and results needed to prove our main theorem. We refer the reader to [6] or [15] for additional background on combinatorial commutative algebra, and to [1] for topological combinatorics. We continue to use the notation from the introduction.
Let $\Delta$ be a simplicial complex on vertex set $V$, which we typically take as $V=[n]:=\{1,2,\dots,n\}$. An element $F\in\Delta$ is called a face, and a face not properly contained in any other face is called a facet. For each $i\in[n]$ we always assume that $\{i\}$ is a face of $\Delta$, which we call a vertex and denote simply as $i$. We will use the notation $\Delta=\langle F_{1},\ldots,F_{s}\rangle$ to denote the complete list of distinct facets. After fixing a vertex $x\in[n]$, we can define two other simplicial complexes from $\Delta$. Namely, the link of $\{x\}$ is the simplicial complex
${\rm link}_{\Delta}(\{x\})=\{G\in\Delta~{}|~{}x\notin G,~{}~{}G\cup\{x\}\in\Delta\},$
@@
${\rm del}_{\Delta}(\{x\})=\{G\in\Delta~{}|~{}x\notin G\}.$
We will abuse notation and write ${\rm link}_{\Delta}(x)$ (respectively, ${\rm del}_{\Delta}(x)$) instead of ${\rm link}_{\Delta}(\{x\})$ (respectively, ${\rm del}_{\Delta}(\{x\})$). A vertex $x$ is a shedding vertex if for any facet $F\in\Delta$ with $x\in F$, there exists a vertex $y\not\in F$ such that $(F\setminus\{x\})\cup\{y\}$ is also a face of $\Delta$. Equivalently, $x$ is a shedding vertex if every facet of ${\rm del}_{\Delta}(x)$ is a facet of $\Delta$.
@@
    $\Delta=\{\emptyset\}$, or $\Delta$ is a simplex (i.e., $\Delta=\langle F\rangle$ for some $F\subseteq[n]$), or
@@
We next collect some results regarding vertex decomposable complexes that will be useful for our study. Recall that a simplicial complex $\Delta$ is a cone if there exists a vertex $v\in V$ such that $v\in F$ for every facet $F\in\Delta$. This implies that there exists a simplicial complex $\Gamma$ on vertex set $V\backslash\{v\}$ such that every face of $\Delta$ is either a face of $\Gamma$, or else of the form $F\cup\{v\}$ for some face $F\in\Gamma$. In this case we write $\Delta=\Gamma\ast v$. In the statement below, the $k$-skeleton of a pure simplicial complex $\Delta$ is (for $k\leq\dim\Delta$) the simplicial complex whose facets are the faces of $\Delta$ having dimension $k$.
@@
    [12, Proposition 3.1.1] All 00-dimensional complexes are vertex decomposable.
@@
    [12, Proposition 3.1.2] A pure $1$-dimensional complex $\Delta$ is vertex decomposable if and only if $\Delta$ is connected.
@@
    [16, Lemma 3.10] Any $k$-skeleton of a vertex decomposable complex $\Delta$ is vertex decomposable.
@@
    [12, Proposition 2.4] A cone $\Delta=\Gamma\ast v$ is vertex decomposable if and only if $\Gamma$ is vertex decomposable.
@@
The following lemma is immediate from the recursive definition of vertex decomposability, and we will use it without comment.
@@
If $\mathcal{F}$ is a family of simplicial complexes, so that every $\Delta\in\mathcal{F}$ is either a simplex or has a shedding vertex v with the property that both $\operatorname{link}_{\Delta}(v)$ and $\operatorname{del}_{\Delta}(v)$ are in $\mathcal{F}$, then every complex in $\mathcal{F}$ is vertex decomposable.
@@ 3. Proof of Main Result
In this section we provide the proof of Theorem 1.1. In what follows, we let $\Delta$ be a pure simplicial complex on $n$ vertices with codimension $c=n-\dim\Delta-1$. If $F$ is a set of $\dim\Delta+1$ vertices that is not a face of $\Delta$, we say $F$ is an antifacet. We write $e(\Delta)$ for the number of facets of $\Delta$ and $\bar{e}(\Delta)$ for the number of antifacets of $\Delta$. Note that $\bar{e}(\Delta)=\binom{n}{c}-e(\Delta)$.
@@
If $\dim\Delta\geq 1$, $c\geq 2$ and $\bar{e}(\Delta)\leq 2c-1$, then $\Delta$ has at least two vertices that each avoids at most $32c-$ 3 antifacets.
@@
Counting by antifacets, we see that
(3.1) $\left|\mathcal{P}\right|\leq c\cdot(2c-1)=2c^{2}-c.$
Suppose for a contradiction that all but at most one vertex avoids more than $32c-$ 3 antifacets. Then
@@
As $22c^{2}-c\geq 2c^{2}-$ 2 holds only if $c\leq 2$, we have reduced to the case $c=2$. But in this case, the inequalities (3.1) and (3.2) give us that $2\cdot 3\geq(n-1)\cdot 2$, so we are left with the case $n=4$.
Thus, the proof reduces to the case where $\Delta$ is a one-dimensional complex on $4$ vertices. In this case, $\bar{e}(\Delta)\leq 3$ and $c=2$, and so we need to show there must be at least two vertices that each avoids at most $12c-3=$ 1 antifacets. Viewing the antifacets as edges in a graph on four vertices, the antifacets give a graph on four vertices where every vertex has degree at most two (since $\Delta$ cannot have an isolated vertex). We finish the proof with the following easily-verified observation: For any graph on four vertices with three or fewer edges, where the degree of each vertex is $\leq 2$, there are at least two vertices that are each disjoint from at most one edge. ∎
@@
We make some other easy observations. Recall that a minimal nonface of $\Delta$ is a nonface of $\Delta$ that is minimal under inclusion with respect to this property. Note that if $\{v,w\}$ is a minimal nonface of $\Delta$ of cardinality two, then $w$ is not a vertex of $\operatorname{link}_{\Delta}(v)$. The following lemma says that minimal nonfaces of cardinality two are rare when $\dim\Delta\neq 1$ and the number of antifacets is small enough (equivalently, if the number of facets is large enough relative to $n$).
@@
If $\bar{e}(\Delta)\leq 2c$, then no nonface of $\Delta$ has fewer than $\dim\Delta$ vertices, and at most one (minimal) nonface has $\dim\Delta$ vertices.
@@
A nonface of size $\dim\Delta-1$ is contained in $\binom{c+2}{2}$ antifacets. But $(c+2)(c+1)/2$ is greater than $2c$ for all $c$, so no such nonface can exist.
Each nonface with $\dim\Delta$ vertices is contained in $c+1$ antifacets. If there existed a pair of such nonfaces, they would be contained in at most one common antifacet. But $12(c+1)-$ 1 is greater than $2c$, contradicting our assumptions. ∎
@@
Suppose $v$ is a vertex of $\Delta$. If $\bar{e}(\Delta)\leq 2c-1$ and $b$ is the codimension of $\operatorname{link}_{\Delta}(v)$, then $\bar{e}(\operatorname{link}_{\Delta}(v))\leq 2b-1$.
@@
Note that the statement is trivial if $\dim\Delta\leq 1$ since in this case $\bar{e}(\operatorname{link}_{\Delta}(v))=0$. Now since $\Delta$ is pure, we have that $\dim\operatorname{link}_{\Delta}(v)=\dim\Delta-1$. If $\dim\Delta>2$, then by Lemma 3.3 we have that $\Delta$ has no nonface of size two, and hence $\operatorname{link}_{\Delta}(v)$ has $n-1$ vertices by the comment after Remark 3.2. From this we get that $b=(n-1)-(\dim\Delta-1)-1=c$. The result then follows by noting that if $F$ is an antifacet of $\operatorname{link}_{\Delta}(v)$, then $F\cup\{v\}$ is an antifacet of $\Delta$.
If $\dim\Delta=2$, then there are two cases. If $\Delta$ has no nonface of size two, then the argument follows as above. Otherwise, $\Delta$ has a nonface of size two, and then again by Lemma 3.3 we have that $b=c-1$ (and hence also $c\geq 1$), and there is a unique minimal nonface $\{v,w\}$ having two vertices. Now, as before there are $\bar{e}(\operatorname{link}_{\Delta}(v))$ antifacets of $\Delta$ containing $v$ but not $w$, and an additional $n-2$ antifacets of $\Delta$ containing $\{v,w\}$. We obtain that $\bar{e}(\operatorname{link}_{\Delta}(v))\leq\bar{e}(\Delta)-(n-2)\leq(2c-1)-(c+1)=2b-c$, which yields the desired result. ∎
@@
With this we have the following observation.
@@
    Any vertex that is contained in at least $c$ antifacets is a shedding vertex.
@@
Recall that $\Delta$ is assumed to be a pure simplicial complex on vertex set $[n]$ with codimension $c$, with at least $\binom{n}{c}-2c+1$ facets, so that $\bar{e}(\Delta)\leq 2c-1$. The statement can be checked directly for all simplicial complexes on $n\leq 3$ vertices, so we suppose that $n>3$. It is convenient to handle the cases when $c=0,1,n-2,n-1$ separately. If $c=0$ or $1$, then $\Delta$ is a simplex, or a simplex boundary with at most one face removed, and the result follows by Lemma 2.4. On the other hand, if $c=n-1$, then $\dim\Delta=0$, and the result is immediate.
Moving forward, we can assume that $c\leq n-2$ (or equivalently, that $\dim\Delta\geq 1$) and that $c\geq 2$. To prove the theorem we wish to find a shedding vertex $v$ so that $\operatorname{link}_{\Delta}(v)$ and $\operatorname{del}_{\Delta}(v)$ also satisfy the hypotheses of the theorem. Note that we are using Lemma 2.5. First observe that for any vertex $v$, the complex ${\rm link}_{\Delta}(v)$ is pure of dimension $\dim\Delta-1$, and also satisfies the hypotheses in Corollary 3.4.
We next turn to the deletion. For this note that if $v$ is a shedding vertex, then by definition $\dim\operatorname{del}_{\Delta}(v)=\dim\Delta$. Since the number of vertices of $\operatorname{del}_{\Delta}(v)$ is $n-1$, we have that the codimension of $\operatorname{del}_{\Delta}(v)$ is $c-1$. Hence to complete the proof we need to find a shedding vertex $v$ satisfying $\bar{e}(\operatorname{del}_{\Delta}(v))\leq 2c-3$.
If every (or even all but one) vertex of $\Delta$ is a shedding vertex, then we are done by Lemma 3.1. Otherwise, Lemma 3.6 yields a shedding vertex $v$ that is contained in at least $c$ antifacets, so that $\bar{e}(\operatorname{del}_{\Delta}(v))\leq 2c-1-c$. In either case, we have produced the desired shedding vertex and the result follows. ∎